OpenStudy (anonymous):

halp xc Angelica's service club is selling flowers for a fundraiser. Each flower costs them $0.50 to purchase. They set the sale price of each flower at the purchase price plus 40%. Write and solve an equation to determine how many flowers they will need to sell to make a profit of $500. Use complete sentences to explain what each number and variable in your equation represents. Write your solution in the context of the original problem.

OpenStudy (anonymous):

it costs them $0.50 to buy each flower they then turn around and sell each flower at $0.50 + 40% of $0.50 = 0.50 + 0.4*0.5 = 0.70 dollars so if they buy x flowers it will cost them a total of 0.50x dollars let's make C the cost function, so C = 0.50x and if they sell all x flowers bought at $0.70, then the revenue function is R = 0.70x the profit, call it P, is equal to the difference of the revenue and cost Profit = Revenue - Cost P = R - C P = 0.70x - 0.50x P = (0.70 - 0.50)x P = (0.20)x P = 0.20x so if you buy and sell x flowers, then your profit is 0.20x dollars you want the profit to be $500, which means P = 0.20x 500 = 0.20x ... plug in P = 500 then solve for x I'll let you do this

OpenStudy (anonymous):

Ohh ok thankyou!

OpenStudy (anonymous):

Anytime :)